Find the number of sides of a polygon if the differences between any two consecutive interior angles of the polygon is 5(degree) and the smallest angle is 120(degree).
Answer by nycsharkman(136)

180(n-2) = n/2( 120 + 120 + 5(n-1))
If you simplify, you get
5n^2 -125n +720 =0
n^2 -25n +144 =0
(n -16)(n-9) =0
n =16 or n =9